You expect the stock market to increase, but instead of acquiring stock, you decide to acquire a stock index futures contract. That index is currently 60.0, and the contract has a value that is $800 times the amount of the index. The margin requirement is $4,000.

  1. When you make the contract, how much must you put up? Round your answer to the nearest dollar.

    $ ________

  2. What is the value of the contract based on the index? Round your answer to the nearest dollar.

    $ ________

  3. If the value of the index rises 1 percent to 60.600, what is the profit on the investment? Round your answer to the nearest cent.

    $ ________

    What is the percentage earned on the funds you put up? Round your answer to one decimal place.

    ________%

  4. If the value of the index declines 1 percent to 59.400, what percentage of your funds will you lose? Round your answer to one decimal place. Enter your answer as a positive value.

    _______ %

  5. What is the percentage you earn (or lose) if the index falls to 55.0? Round your answer to one decimal place. Enter your answer as a positive value.

    The percentage __profit / loss___ is ______%.
Answer #1

Ans. a) Given Index Value 60.00, Lot Size 800 and Margin required $4000

The value need to put up to acquire Index future is the amount of margin required.

The value need to put up to acquire Index future = $4000

Ans. b) The Value of the contract = Index value * Lot Size

  The Value of the contract = 60*800

  The Value of the contract = $48000

Ans. c): If Index rises 1% from 60.00 to 60.600

Profit = Increment in Index value * Lot size

Profit = (60.600-60) * 800

Profit = 0.600*800 = $480

The Fund Invested = Margin paid = $4000

Percentage Earned on the fund invested = (Profit/ Fund Invested)*100

Percentage Earned on the fund invested = 480/4000 = 12.00%

Ans.d) If Index declines 1% from 60.00 to 59.400

Loss = Decrease in Index value * Lot size

Loss = 0.600*800 = $480.00

Percentage lose on the fund invested = (Loss/Fund Invested)*100

Percentage lose on the fund invested = 480/4000 = 12%